OpenStudy (jchick):

Someone please help me Which of the following equations is written in slope-intercept form? Choose one answer. a. 2x + 3y = 6 b. x = -three halvesy + 3 c. 3y = -2x + 6 d. y = -two thirdsx + 2

OpenStudy (jchick):

I don't understand this

OpenStudy (anonymous):

a quick google search tells us that slope intercept form is "y = mx + b" do you see something that matches that?

OpenStudy (jchick):

@undeadknight26 can you help me here?

OpenStudy (anonymous):

@jchick slope intercept form is: y = mx + b

OpenStudy (jchick):

Yes I got that but I don't really understand

OpenStudy (jchick):

I am not very good at this

OpenStudy (anonymous):

ok do you want to understand? or just want the answer?

OpenStudy (jchick):

I want to understand it

OpenStudy (jchick):

I don't just want the answer I need to learn it for future reference

OpenStudy (jchick):

However I tried to come up with an answer do you want to see it?

OpenStudy (jchick):

I think the answer is A

OpenStudy (anonymous):

the answer is not A

OpenStudy (jchick):

Ok

OpenStudy (jchick):

Can you help me to understand then?

OpenStudy (anonymous):

ok i will tell you something which will basically give you the answer, but only if you promise to stay and learn the whole thing, deal? it will take 15 min or so

OpenStudy (jchick):

Ok that is fine thanks btw

OpenStudy (jchick):

I really want to understand this

OpenStudy (anonymous):

ok so again slope intercept form is y = mx + b

OpenStudy (anonymous):

so looking at your possible answers, do you see something that has a "y =" on the left side? just like the slope intercept form?

OpenStudy (jchick):

Yes I just saw that

OpenStudy (jchick):

Ok so still how do you figure it out if it was harder?

OpenStudy (jchick):

Because I don't think that they will all be like that

OpenStudy (anonymous):

ok so this is the equation for a line, y = mx + b

OpenStudy (jchick):

Ok

OpenStudy (anonymous):

the y is our output, and the x is our input, the m is our slope, and the b is our y intercept, we can go over those in detail if you like

OpenStudy (jchick):

Ok please do

OpenStudy (jchick):

thanks for helping me btw

OpenStudy (anonymous):

you are welcome

OpenStudy (anonymous):

ok so let me draw the line first, or can you draw it?

OpenStudy (jchick):

I am not sure how to

OpenStudy (jchick):

I am still new to this site

OpenStudy (anonymous):

you click on the blue draw button

OpenStudy (jchick):

Oh ok

OpenStudy (jchick):

How do you want the line drawn?

OpenStudy (jchick):

Like where?

OpenStudy (anonymous):

ok good, so lets start with the information we have, so we have y = -2/3x + 2, so can you tell me a point i can plot right away just by looking at that?

OpenStudy (anonymous):

@jchick , i will tell you a point i can plot right away looking at y = -2/3x + 2, is (0, 2)

OpenStudy (jchick):

Sorry i misunderstood what you were asking

OpenStudy (anonymous):

y = -2/3x + 2 match that to y = mx + b y = y m = -2/3 x = x b = 2

OpenStudy (anonymous):

do you see how those match up?

OpenStudy (jchick):

Yes I do now

OpenStudy (anonymous):

ok pefect, so what is the value of b for our equation?

OpenStudy (jchick):

2 right?

OpenStudy (anonymous):

yep b is 2, pefect, b is called the y intercept, so 2 is our y-intercept, that means that the line intercepts the y axis when y = 2 ,

OpenStudy (anonymous):

and we know the y axis is when x = 0, so we have the point (0,2), follow me?

OpenStudy (jchick):

Ok awesome I get it now!!

OpenStudy (anonymous):